40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Sep 8 08:24:49 2014, Steve, Update, Computer Scripts / Programs, preparing vac system to reboot 
    Reply  Mon Sep 8 11:10:26 2014, ericq, Update, Computer Scripts / Programs, preparing vac system to reboot 
       Reply  Mon Sep 8 11:34:47 2014, rana, Update, Computer Scripts / Programs, preparing vac system to reboot 
          Reply  Mon Sep 8 12:11:36 2014, manasa, Update, Computer Scripts / Programs, preparing vac system to reboot 
             Reply  Mon Sep 8 15:50:47 2014, ericq, Update, Computer Scripts / Programs, +15V PSL Sorensen replaced 
       Reply  Tue Sep 9 09:19:21 2014, Steve, Update, safety, response to fried Sorensen friedSorenson.pngsmokeAlarm1a.jpgsmokeAlarm2a.jpg
       Reply  Tue Sep 9 14:18:40 2014, Steve, Update, Computer Scripts / Programs, preparing vac system to reboot v1closedfor1h.png
       Reply  Tue Sep 16 17:02:41 2014, Steve, Update, Computer Scripts / Programs, PSL output shutter is floating 
Message ID: 10477     Entry time: Tue Sep 9 14:18:40 2014     In reply to: 10468
Author: Steve 
Type: Update 
Category: Computer Scripts / Programs 
Subject: preparing vac system to reboot 

Quote:

Quote:

Q and Steve will follow elog 10028 entry to prepare the vacuum system for safe reboot

Here's the sequence of the morning so far:

  • I aligned the IFO (IR arms with ASS, X green with PZTs, PRM with PRMI locked on REFL33)
  • I closed the PSL shutter, and went inside to align PRM and both ITM oplevs (all others were within 10urad of zero in both directions)
  • While aligning those oplevs, I noticed the smell of burnt electronics. We tracked it down to the +15V sorensen in the rack nearest the PSL table
    • I claim the precipitating event was PSL shutter activity. If I recall correctly, the seismic rainbow traces went bonkers around the same time as the shutter was closed. There is a Guralp interface in the rack powered by the failed sorensen, so this would explain the erratic seismometer signals correlated with the power supply failure. We will look into potential shorts caused by the shutter. (Steve looked up the PMC trans and Guralp DQ channels, and confirmed the temporal coincidence of the events.)
  • We shut off all of the sorensens so that electronics were not being driven asymmetrically. 
  • Steve and I secured the vacuum system for computer reboots, as referred to in Steve's elog. Some combination of Jenne, Rana and Manasa shut down the control room computers, and turned off the watchdogs. 
  • Manasa and I moved Chiara inside, next to Mafalda, along with its backup HDs. It has been labeled. 
  • Booted up control room machines, they came up happy. 
  • FB and front-ends didn't need reboot, for some lucky reason. Watchdogs came back happily, oplev spots didn't move noticeably. 

The IFO is still down, as the PMC won't lock without the rack power, and we haven't pinned down the shorting mechanism. We don't want the replacement sorensen to immediately blow when plugged in. 

Vacuum safe reboot required one hour of no pumping of the vac envelope.

Attachment 1: v1closedfor1h.png  73 kB  | Hide | Hide all
v1closedfor1h.png
ELOG V3.1.3-